OpenStudy (anonymous):

Y=x^2+5x+6 graph each parabola, give its vertex axis, x-intercept,y intercept. Can someone teach me how to do this?

OpenStudy (anonymous):

k, let's walk through it.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

See if this makes sense: https://drive.google.com/file/d/0B5q8dbsgquFmQjJFcXJjVkgySFE/edit?usp=sharing

OpenStudy (anonymous):

I cant see anything

OpenStudy (anonymous):

Nevermind got it...hold on

OpenStudy (anonymous):

Okay i understand....Next

OpenStudy (anonymous):

So the x intercept is -2,-3

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There is two of them in this case.

OpenStudy (anonymous):

okay....Next

OpenStudy (anonymous):

The graph crosses the x-axis twice: https://www.desmos.com/calculator/xx0vrdccd8

OpenStudy (anonymous):

You get how to use the quadratic formula and plug in "A", "B", and "C"?

OpenStudy (anonymous):

I have the answers for these questions and they dont match what you did

OpenStudy (anonymous):

answer wise or process wise?

OpenStudy (anonymous):

The Vertex is (-5/4,-1/4); axis is x=-5/2

OpenStudy (anonymous):

yes, we haven't done the vertex yet.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

This part may be confusing if you're not good at fractions

OpenStudy (anonymous):

you can zoom into the picture.

OpenStudy (anonymous):

@jbell0385 how did you got 24/4?

OpenStudy (anonymous):

I wanted all the bottoms to be the same number (4). 24/4 equals 6, so they are the same thing.

OpenStudy (anonymous):

Once you have the same bottom number, you can add the top numbers.

OpenStudy (anonymous):

I get it..So we got the vertex...now why does it say the axis is x=-5/2??

OpenStudy (anonymous):

It's a little hard to explain over the internet, but the vertex is the very bottom point of the parabola.

OpenStudy (anonymous):

It's basically in the middle of the shape.

OpenStudy (anonymous):

i know what your talking about

OpenStudy (anonymous):

So if you drew a line straight up and down at x=-5/2 you would chop the parabola in half.

OpenStudy (anonymous):

So they call it the line of symmetry. It will always be the "x" value of your vertex.

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

You just have to get used to doing those 3 steps over and over until it becomes comfortable.